Simplify. -2|3-2|+6-12|-2(3+8-2^3)| . Thanks for your help.
Answer by Alan3354(69443)

-2|3-2|+6-12|-2(3+8-2^3)|
= -2|1|+6-12|-2(3+8-8)|
= -2 + 6 -12|-2(3)|
= 4 -12|-6|
= 4 -12*6
= 4 - 72
= -68